The volume of a 10-ounce box of cereal is 182.743 cubic inches. The width of the box is 3.1875 inches less than the length, and its depth is 2.31245 inches. Find the length and width of the box to the nearest thousandths.
Answer by ankor@dixie-net.com(22740)

The volume of a 10-ounce box of cereal is 182.743 cubic inches. The width of the box is 3.1875 inches less than the length, and its depth is 2.31245 inches. Find the length and width of the box to the nearest thousandths.
:
The one who chose the numbers in this problem should get back on their medication!
:
Let x = the length of the box
Then (x-3.1875) = the width
:
A simple volume equation:
x * (x-3.1875) * 2.31245 = 182.743
:
(x^2 - 3.1875x) * 2.31245 = 182.743
:
2.31245x^2 - 7.370934375x - 182.743 = 0;
a quadratic equation which has to be solved using the quadratic formula:
a = 2.31245; b = 7.370934375; c = 182.743
:
Do this and you should get (positive solution is what we want here):
x = +10.625126 is the length
:
10.625126 - 3.1875 = 7.437626 is the width
:
Check our solution:
10.625126 * 7.437626 * 2.31245 = 182.743
:
Is this some kind of punishment? Hope you don't have very many like this.